# FALCON: Fairness Learning via Contrastive Attention Approach to Continual Semantic Scene Understanding in Open World

### v1

@@ Abstract
Continual Learning in semantic scene segmentation aims to continually learn new unseen classes in dynamic environments while maintaining previously learned knowledge. Prior studies focused on modeling the catastrophic forgetting and background shift challenges in continual learning. However, fairness, another major challenge that causes unfair predictions leading to low performance among major and minor classes, still needs to be well addressed. In addition, prior methods have yet to model the unknown classes well, thus resulting in producing non-discriminative features among unknown classes. This paper presents a novel Fairness Learning via Contrastive Attention Approach to continual learning in semantic scene understanding. In particular, we first introduce a new Fairness Contrastive Clustering loss to address the problems of catastrophic forgetting and fairness. Then, we propose an attention-based visual grammar approach to effectively model the background shift problem and unknown classes, producing better feature representations for different unknown classes. Through our experiments, our proposed approach achieves State-of-the-Art (SOTA) performance on different continual learning settings of three standard benchmarks, i.e., ADE20K, Cityscapes, and Pascal VOC. It promotes the fairness of the continual semantic segmentation model.
@@ 1 Introduction
The semantic segmentation networks, e.g., Transformers [56] and Convolutional Neural Networks [7], learned from data with a closed-set of known classes have shown outstanding performance. However, they often suffer performance degradation when encountering novel objects or classes in new dynamic environments [19, 3, 52]. To improve their performance, several transfer learning and domain adaptation methods [26, 1, 52] were introduced to adapt trained models into deployed environments. While the former often aims to fine-tune the model on labeled data collected in the new environments, the latter adapts the model to the new domains in an unsupervised manner [1, 26]. However, these methods cannot handle novel objects well due to their close-set learning. In practice, the semantic segmentation models should be able to adaptively and continually learn the new knowledge of novel classes. It motivates the development of Continual Learning paradigm [3, 19, 40], a.k.a, Continual Semantic Segmentation (CSS), where the segmentation models are learned sequentially to new contents of data.
@@
Far apart from prior segmentation methods [7, 56] that learn one time on static, closed-set data, Continual Learning requires the segmentation models to learn from dynamic, open-set data [19, 52]. In addition, in particular scenarios, accessing previous learning data is restricted due to privacy concerns. In CSS, three challenges have been identified, including (1) Catastrophic Forgetting, (2) Background Shift, and (3) Fairness. While the catastrophic forgetting problem [44, 23, 50] depicts the segmentation model tends to forget its knowledge when learning new data, background shift indicates the problem of classes of previous or future data (unknown classes) have collapsed into a background class [3, 19, 60]. Prior methods [19, 40, 5, 52] addressed these two problems by introducing knowledge distillation and pseudo labels. However, these methods can not handle unknown classes since they either consider these unknown classes as a background class or assign unknown pixels by a pseudo label of prior known classes [19, 52]. More importantly, the last problem, fairness, is a significant challenge that limits the performance of CSS models.
As shown in Fig. 2, the number of pixels of each class in training data have been imbalanced among classes and significantly decreased after each task. Thus, this bias influences the learning procedure and model predictions that later cause unfair predictions among classes. However, limited studies are taking the fairness problem into account. [51] presented a similar problem in domain adaptation and extended it to continual learning [52]. These methods rely on the assumption of ideal fair or balanced data distributions. However, it is not applicable in practice since the size, i.e., the number of pixels, of several classes can never be more significant than others. For example, the size of the bottle should not be more significant than the size of a car. Meanwhile, the current knowledge distillation methods [19, 39, 4] in CSS are unable to handle the fairness problem since they focus on modeling catastrophic forgetting and background shift problems. Therefore, it is essential to develop a new CSS approach to address these limitations.
Contributions of This Work: This work presents a novel Fairness Learning via Contrastive Attention Approach (FALCON) to Continual Semantic Segmentation (as shown in Fig. 1). First, we introduce a novel Contrastive Clustering Paradigm approach to Continual Learning that models the catastrophic forgetting problem. Second, by analyzing the limitation of vanilla Contrastive Clustering in biased data, we introduce a novel Fairness Contrastive Clustering loss to model the fairness problem in continual learning efficiently. Third, to effectively model the background shift problem, we introduce a new Attention-based Visual Grammar that model the topological structures of feature distribution to handle the unknown classes effectively. Finally, the ablation studies illustrate the effectiveness of the proposed approach in different aspects of fairness promotion in CSS models. Compared with prior methods, our approach achieves state-of-the-art performance on different settings of three standard benchmarks of CSS, including ADE20K, Pascal VOC, and Cityscapes.
@@ 2 Related Work
Continual Semantic Segmentation Several studies have been introduced to address catastrophic forgetting and background shift problems in CSS [20, 28, 32, 18, 19, 38, 37, 34]. The common approach of CSS adopts knowledge distillation [19] and pseudo labels [3] to model catastrophic forgetting and background shift, respectively. Later, it is further improved by decoupling knowledge representations [60], modeling the inter- and intra-class knowledge [40], distinguishing the feature representations of the future classes [5, 39], reducing background confusion [59], or modeling distillation loss via the geodesic flow [48]. Another approach [4] adopted the mask-based segmentation networks [12, 13] to improve the performance of CSS models. Recent studies have introduced CSS under the unsupervised domain adaptation settings [54, 45, 46]. However, prior studies have yet well modeled the unknown classes in the open-world environments. Particularly, as previous methods [19, 4] use the pseudo labels to model the unknown classes, the future classes will be treated as a background class. Then, Joseph et al. [27] improves the unknown class modeling by using clustering but this method considers all different unknown classes as a single cluster, leading to the non-discriminative features among unknown classes.
@@
Contrastive Learning is a common learning approach [36, 10, 9] to structure the deep feature representations in the deep latent space. Oorde et al. [36] first introduced the Noise-Contrastive Estimation (InfoNCE) learning framework. Then, Chen et al. [9] presented SimCLR, a self-supervised contrastive learning approach to improve the representation power of Residual Networks. He [25] proposed a Momentum Contrast framework for unsupervised representation learning. Later, it was further improved by using MLP projection head [10] and extended to improve the self-supervised training process of vision transformers [11]. Cui et al. [16] introduced a supervised parametric contrastive learning loss to address the long-tailed recognition. Li et al. [30] adopted contrastive learning to develop the one-stage online contrastive clustering method. Radford et al. [41] presents a contrastive framework to learn the vision-language model. Later, several methods also adopted this framework to vision-language pretraining [41, 29].
@@
Imbalanced and Fairness Learning The early methods utilized the balanced Softmax loss [43] to alleviate the impact of imbalanced data distribution. Later, Wang et al. [55] introduced a Seesaw loss to re-balance the contributions of positive and negative instances via the mitigation and compensation modules. Ziwei et al. [31] introduced a dynamic meta-embedding to model the imbalanced classification problem. Chu et al. [14] reduce the bias in the segmentation model by presenting a new stochastic training scheme. Szabo et al. [49] presented a tilted cross-entropy loss to promote class-relevant fairness. However, there are limited studies that address the fairness problem in CSS. Truong et al. [51] introduced a fairness domain adaptation approach to semantic segmentation and later extended it into continual learning setting [52]. However, these methods [51, 52] rely on the assumption of ideal balanced data which could not be achieved by nature. To address the limitations in prior work, this paper will introduce a novel approach to effectively model the fairness problem and unknown classes in the continual learning setting.
Figure 3: The Proposed Open-World Fairness Continual Learning Framework.
@@ 3.1 Continual Learning via Contrastive Clustering
Apart from prior methods [19, 4, 39], our CSS is defined as Contrastive Clustering Learning. Given a set of centroid vectors $\{\mathbf{c}_{i}\}_{i=1}^{N_{K}+N_{U}}$ where $N_{K}=|\mathcal{C}^{1..t}|$ and $N_{U}$ is the number of known and unknown classes up to current learning tasks. Prior work [27, 52, 4] often defined the number of unknown classes as $1$ where background classes are considered as a single unknown class. Formally, our Contrastive Clustering Learning for CSS can be defined as Eqn. (2).
@@
By defining CSS as contrastive clustering learning, the knowledge of the segmentation model has been well maintained via the cluster vectors $\mathbf{c}$. Then, minimizing Eqn. (2) will separate the representations of different classes while gathering the features of the same class into the same cluster. As the cluster vectors $\mathbf{c}$ of the old classes $\mathcal{C}^{1..t-1}$ have been well learned to represent their knowledge, these vectors are frozen at learning step $t$ to maintain the knowledge representations of previous classes to address the catastrophic forgetting problem. To effectively learn cluster vectors $\mathbf{c}$, the cluster vector $\mathbf{c}$ will periodically updated after each $M$ steps by the momentum update [27, 52, 25] based on the features $\mathbf{f}^{t}_{h,w}$ assigned to cluster $\mathbf{c}$. However, there are two major problems in contrastive clustering learning. First, since the training data in CSS suffer the bias among classes as shown in Fig. 2, this bias will influence Eqn. (2) and cause the unfair predictions. Second, as the function $\phi$ requires the labels to determine the features belonging to clusters, it limits the ability to model the unknown classes where the labels are not available. Therefore, Secs. 3.2-3.3 will present a novel approach to tackle these problems.
@@ 3.2 Fairness Contrastive Clustering Learning
While the contrastive clustering learning defined in Eqn. (2) promotes the compact representations of features around their clusters, inspired by [62, 16, 8], we observe that the imbalanced class distribution will influence unfair behaviors among classes. In particular, for simplicity, we consider $\{\mathbf{f}^{t}_{i}\}_{i=1}^{L}$ is the set of features that belong to the cluster $\mathbf{c}$ at learning step $t$ (i.e., $\phi(\mathbf{f}_{i}^{t},\mathbf{c})=1$) and $L$ is the number of features (in this case, $L$ is the total number of pixels belong to the class of cluster $\mathbf{c}$). Let us define the enforcement between the feature $\mathbf{f}^{t}_{t}$ and the cluster $\mathbf{c}$ as $\ell_{i}=\frac{\exp(\mathbf{f}^{t}_{i}\times\mathbf{c})}{\sum_{\mathbf{f}^{\prime}}\exp(\mathbf{f}^{\prime}\times\mathbf{c})}$. Hence, the lower the value of the enforcement $\ell_{i}$ is, the more compact the representation of visual features and clusters is. Then, the contrastive clustering learning loss in Eqn. (2) of the entire cluster $\mathbf{c}$ can be defined as Eqn. (3).
@@
Proposition 1: If the contrastive clustering loss $\mathcal{L}_{Cont}(;,\mathbf{c})$ achieves the optimal value, the enforcement $\ell_{i}$ between the feature and the cluster will converge to $\ell_{i}=L^{-1}$.
Proposition 1 has implied that the class with more samples will result in a lower value of the enforcement and produce a more compact representation while the class having fewer samples will be more scattered in the feature space due to the higher value of the enforcement. In particular, let $L_{major}$ and $L_{minor}$ be the number of samples of the major and minor class where $L_{major}>>L_{minor}$. Then, based on Proposition 1, the enforcement between features and the cluster of the major class will be significantly lower than the one of the minor class, i.e., $L_{major}^{-1}<<L_{minor}^{-1}$. Therefore, a direct adoption of the contrastive clustering loss in Eqn. (2) will result in an unfair CSS model. In addition, for classes in the minority group, the weak enforcement results in the feature presentations of classes being far away from their clusters. Thus, the model will produce non-discriminative features compared to the ones in the majority group. Moreover, if the loss is applied to the cases of unknown labels, these feature representations can be scattered in the latent space and pulled into the incorrect clusters due to weak enforcement between features and clusters (Fig. 4).
To address the unfair problem in contrastive clustering learning, inspired by [62, 16, 8], we introduce a scaling factor $\alpha$ and a learnable transition vector $\mathbf{v}$ for each cluster $\mathbf{c}$ (all clusters have the same value of $\alpha$ but different vector $\mathbf{v}$). Our Fairness Contrastive Clustering Learning Loss for the entire cluster in Eqn. (3) can be re-formed as:
@@
Proposition 2: If the fairness contrastive clustering loss $\mathcal{L}^{\alpha}_{Cont}(;,\mathbf{c})$ achieves the optimal value, the enforcement $\ell_{i}$ between the feature and the cluster will converge to $\ell_{i}=(\alpha^{-1}+L)^{-1}$. Proofs of Propositions 1-2 are provided in the supplementary.
@@
3.3 Open-world Unknown Class Modeling

### v3

# FALCON  [Uncaptioned image]: Fairness Learning via Contrastive Attention Approach to Continual Semantic Scene Understanding

@@ Abstract
Continual Learning in semantic scene segmentation aims to continually learn new unseen classes in dynamic environments while maintaining previously learned knowledge. Prior studies focused on modeling the catastrophic forgetting and background shift challenges in continual learning. However, fairness, another major challenge that causes unfair predictions leading to low performance among major and minor classes, still needs to be well addressed. In addition, prior methods have yet to model the unknown classes well, thus resulting in producing non-discriminative features among unknown classes. This work presents a novel Fairness Learning via Contrastive Attention Approach to continual learning in semantic scene understanding. In particular, we first introduce a new Fairness Contrastive Clustering loss to address the problems of catastrophic forgetting and fairness. Then, we propose an attention-based visual grammar approach to effectively model the background shift problem and unknown classes, producing better feature representations for different unknown classes. Through our experiments, our proposed approach achieves State-of-the-Art (SoTA) performance on different continual learning benchmarks, i.e., ADE20K, Cityscapes, and Pascal VOC. It promotes the fairness of the continual semantic segmentation model.
@@ 1 Introduction
The semantic segmentation networks, e.g., Transformers [70] and Convolutional Neural Networks [8], learned from data with a closed-set of known classes, have shown outstanding performance. However, they often suffer performance degradation when encountering novel objects or classes in new dynamic environments [21, 3, 63, 64]. To improve their performance, several transfer learning and domain adaptation methods [29, 1, 63, 60, 65, 58, 41, 61, 30, 59] were introduced to adapt trained models into deployed environments. While the former often aims to fine-tune the model on labeled data collected in the new environments, the latter adapts the model to the new domains in an unsupervised manner [1, 29]. However, these methods cannot handle novel objects well due to their close-set learning. In practice, the semantic segmentation models should be able to adaptively and continually learn the new knowledge of novel classes. It motivates the development of Continual Learning paradigm [3, 21, 47], a.k.a, Continual Semantic Segmentation (CSS), where the segmentation models are learned sequentially to new contents of data.
@@
Far apart from prior segmentation methods [8, 70] that learn one time on static, closed-set data, Continual Learning requires the segmentation models to learn from dynamic, open-set data [21, 63]. In a particular scenario, accessing previous learning data is restricted due to privacy concerns. In CSS, three challenges have been identified, including (1) Catastrophic Forgetting, (2) Background Shift, and (3) Fairness. While the catastrophic forgetting problem [51, 25, 57] depicts the segmentation model tends to forget its knowledge when learning new data, background shift indicates the problem of classes of previous or future data (unknown classes) have collapsed into a background class [3, 21, 76]. Prior methods [21, 47, 5, 63] addressed these two problems by introducing knowledge distillation and pseudo labels. However, these methods can not handle unknown classes since they either consider these unknown classes as a background class or assign unknown pixels by a pseudo label of prior known classes [21, 63]. More importantly, the last problem, fairness, is a significant challenge that limits the performance of CSS models (Fig. 1).
@@
As shown in Fig. 2, the number of pixels of each class in training data have been imbalanced among classes and significantly decreased after each task. Thus, this bias influences the learning procedure and model predictions that later cause unfair class predictions. However, limited studies are taking the fairness problem into account. [62] presented a similar problem in domain adaptation and extended it to continual learning [63]. These methods rely on the assumption of ideal fair or balanced data distributions. However, it is not applicable in practice since the size, i.e., the number of pixels, of several classes can never be more significant than others. For example, the bottle size should not be more significant than the size of a car. Meanwhile, the current knowledge distillation methods [21, 46, 4] in CSS are unable to handle the fairness problem since they focus on modeling catastrophic forgetting and background shift problems. Therefore, it is essential to develop a new CSS approach to address these limitations.
Contributions of This Work. This work presents a novel Fairness Learning via Contrastive Attention Approach (FALCON) to Continual Semantic Segmentation (as shown in Fig. 1). Our contributions can be summarized as follows. First, we introduce a novel Contrastive Clustering Paradigm approach to Continual Learning that models the catastrophic forgetting problem. Second, by analyzing the limitation of vanilla Contrastive Clustering in biased data, we introduce a novel Fairness Contrastive Clustering loss to model the fairness problem in continual learning efficiently. Third, to effectively model the background shift problem, we introduce a new Attention-based Visual Grammar that model the topological structures of feature distribution to handle the unknown classes effectively. Finally, the ablation studies illustrate the effectiveness of the proposed approach in different aspects of fairness promotion in CSS models. Compared with prior methods, our approach achieves SoTA performance on different settings of three standard benchmarks of CSS, including ADE20K, Pascal VOC, and Cityscapes.
@@ 2 Related Work
Continual Semantic Segmentation. Several studies were introduced to address catastrophic forgetting and background shift problems [22, 33, 38, 20, 21, 45, 44, 40]. The common CSS approach adopts knowledge distillation [21] and pseudo labels [3] to model catastrophic forgetting and background shift, respectively. Later, it was further improved by decoupling knowledge representations [76], modeling the inter- and intra-class knowledge [47, 16], distinguishing the feature representations of the future classes [5, 46, 68, 71], reducing background confusion [74, 75], using visual prompting [32, 36], or modeling distillation loss via the geodesic flow [55]. Another approach [4, 26, 6] adopted the mask-based segmentation networks [13, 14] to improve the performance. Recent studies have introduced CSS under the unsupervised domain adaptation settings [67, 52, 53, 64]. However, prior studies have yet to well model different unknown classes. Particularly, as previous methods [21, 4] use the pseudo labels to model the unknown classes, the future classes will be treated as a single background class. Then, Joseph et al. [31] improves the unknown class modeling using clustering but this method considers all different unknown classes as a single cluster, leading to the non-discriminative features among unknown classes.
@@
Contrastive Learning (CL). CL is a common learning approach [43, 11, 10] to structure the deep feature representations in the deep latent space. Oorde et al. [43] first introduced the Noise-Contrastive Estimation (InfoNCE) learning framework. Then, Chen et al. [10] presented SimCLR, a self-supervised contrastive learning approach to improve the representation power of Residual Networks. He [28] proposed a Momentum Contrast framework for unsupervised representation learning. Later, it was further improved by using MLP projection head [11] and extended to improve the self-supervised training process of vision transformers [12]. Cui et al. [18] introduced a supervised parametric contrastive learning loss to address the long-tailed recognition. Li et al. [35] adopted contrastive learning to develop the one-stage online contrastive clustering method. Radford et al. [48] presents a contrastive framework to learn the vision-language model. Later, several methods also adopted this framework for vision-language pretraining [48, 34].
@@
Imbalanced and Fairness Learning. The early methods utilized the balanced Softmax loss [50] to alleviate the impact of imbalanced data distribution. Later, Wang et al. [69] introduced a Seesaw loss to re-balance the contributions of positive and negative instances via the mitigation and compensation modules. Ziwei et al. [37] introduced a dynamic meta-embedding to model the imbalanced classification problem. Chu et al. [15] reduce the bias in the segmentation model by presenting a new stochastic training scheme. Szabo et al. [56] presented a tilted cross-entropy loss to promote class-relevant fairness. However, there are limited studies that address the fairness problem in CSS. Truong et al. [62] introduced a fairness domain adaptation approach to semantic segmentation and later extended it into continual learning setting [63]. However, these methods [62, 63] rely on the assumption of ideal balanced data, which could not be achieved by nature. To address the limitations in prior work, this paper will introduce a novel approach to effectively model the fairness problem and unknown classes in the continual learning setting.
Figure 3: The Proposed Fairness Learning via Contrastive Attention Approach to Continual Semantic Scene Understanding.
@@ 3.1 Continual Learning via Contrastive Clustering
Apart from prior methods [21, 4, 46], our CSS (Fig. 3) is defined as Contrastive Clustering Learning. Given a set of centroid vectors $\{\mathbf{c}_{i}\}_{i=1}^{N_{K}+N_{U}}$ where $N_{K}=|\mathcal{C}^{1..t}|$ and $N_{U}$ is the number of known and unknown classes up to current learning tasks. Prior work [31, 63, 4] often defined the number of unknown classes as $1$ where background classes are considered as a single unknown class. Formally, our Contrastive Clustering Learning for CSS can be defined as in Eqn. (2).
@@
By defining CSS as contrastive clustering learning, the knowledge of the segmentation model has been well maintained via the cluster vectors $\mathbf{c}$. Then, minimizing Eqn. (2) will separate the representations of different classes while gathering the features of the same class into the same cluster. As the cluster vectors $\mathbf{c}$ of the old classes $\mathcal{C}^{1..t-1}$ have been well learned to represent their knowledge, these vectors are frozen at learning step $t$ to maintain the knowledge representations of previous classes to address the catastrophic forgetting problem. To effectively learn cluster vectors $\mathbf{c}$, the cluster vector $\mathbf{c}$ will periodically updated after each $K$ steps by the momentum update [31, 63, 28] based on the features $\mathbf{f}^{t}_{h,w}$ assigned to cluster $\mathbf{c}$. However, there are two major problems in contrastive clustering learning. First, since the training data suffer the bias among classes as shown in Fig. 2, this bias will influence Eqn. (2) and cause the unfair predictions. Second, as the function $\phi$ requires the labels to determine the features belonging to clusters, it limits the ability to model the unknown classes where the labels are unavailable. Therefore, Secs. 3.2-3.3 will present a novel approach to tackle these problems.
@@ 3.2 Fairness Contrastive Clustering Learning
While contrastive clustering learning in Eqn. (2) promotes the compact representations of features around their clusters, inspired by [78, 18, 9], we observe that the imbalanced class distribution will influence unfair behaviors among classes. In particular, for simplicity, we consider $\{\mathbf{f}^{t}_{i}\}_{i=1}^{L}$ is the set of features that belong to the cluster $\mathbf{c}$ at learning step $t$ (i.e., $\phi(\mathbf{f}_{i}^{t},\mathbf{c})=1$) and $L$ is the number of features (in this case, $L$ is the total number of pixels belong to the class of cluster $\mathbf{c}$). Let us define the enforcement between the feature $\mathbf{f}^{t}_{t}$ and the cluster $\mathbf{c}$ as $\ell_{i}=\frac{\exp(\mathbf{f}^{t}_{i}\times\mathbf{c})}{\sum_{\mathbf{f}^{\prime}}\exp(\mathbf{f}^{\prime}\times\mathbf{c})}$. Hence, the lower the value of the enforcement $\ell_{i}$ is, the more compact the representation of visual features and clusters is. Then, the contrastive clustering learning loss in Eqn. (2) of entire cluster $\mathbf{c}$ can be defined as in Eqn. (3).
@@
Proposition 1. If the contrastive clustering loss $\mathcal{L}_{Cont}(;,\mathbf{c})$ achieves the optimal value, the enforcement $\ell_{i}$ between the feature and the cluster will converge to $\ell_{i}=L^{-1}$.
Proposition 1 has implied that the class with more samples will result in a lower value of the enforcement and produce a more compact representation while the class having fewer samples will be more scattered in the feature space due to the higher value of the enforcement. In particular, let $L_{major}$ and $L_{minor}$ be the number of samples of the major and minor class where $L_{major}>>L_{minor}$. Then, based on Proposition 1, the enforcement between features and the cluster of the major class will be significantly lower than the one of the minor class, i.e., $L_{major}^{-1}<<L_{minor}^{-1}$. Therefore, a direct adoption of the contrastive clustering loss in Eqn. (2) will result in an unfair CSS model. In addition, for classes in the minority group, the lower value of $L$ results in the feature presentations of classes being far away from their clusters. Thus, the model will produce non-discriminative features compared to the ones in the majority group. Moreover, if the loss is applied to the cases of unknown labels, these feature representations can be scattered in the latent space and pulled into the incorrect clusters due to weak enforcement between features and clusters (Fig. 4).
To address the unfair problem in contrastive clustering learning, inspired by [78, 18, 9], we introduce a scaling factor $\alpha$ and a learnable transition vector $\mathbf{v}$ for each cluster $\mathbf{c}$ (all clusters have the same value of $\alpha$ but different vector $\mathbf{v}$). Our Fairness Contrastive Clustering Learning Loss for the entire cluster in Eqn. (3) can be re-formed as follows,
@@
Proposition 2. If the fairness contrastive clustering loss $\mathcal{L}^{\alpha}_{Cont}(;,\mathbf{c})$ achieves the optimal value, the enforcement $\ell_{i}$ between the feature and the cluster will converge to $\ell_{i}=(\alpha^{-1}+L)^{-1}$. Proofs of Propositions 1-2 are in the appendix.
Figure 4: The Enforcement Loss of Contrastive Clustering $\mathcal{L}_{Cont}$ and Fairness Contrastive Clustering $\mathcal{L}^{\alpha}_{Cont}$ on Pascal VOC. Since $\mathcal{L}_{Cont}$ suffers severe biased, its clusters of minor classes remain scattered. Our $\mathcal{L}^{\alpha}_{Cont}$ produces a more uniform loss among classes, promotes fairness and compactness of clusters.
@@
3.3 An Efficient Unknown Class Modeling